# envname-P envname#1

### v2

@@ Appendix E Training Curves
Figure 6 presents the validation curves during end-to-end training on the NQ and Musique datasets. Recall scores consistently increase while evaluation losses steadily decrease, indicating stable and effective optimization. Higher compression ratios generally yield lower recall and higher loss, mirroring the trends observed during the compression pretraining stage.
@@ F.1 Effect of Freezing the Compressor and Query Reasoner
We investigate the effect of limiting the fine-tuning scope to the generator module while freezing both the compressor and query reasoner. Specifically, we examine two representative compression settings, CR=32 and CR=128, and compare model performance when only the generator is fine-tuned during both the instruction tuning and end-to-end QA training stages. The results are shown in Table 13 and 14. During the compression learning and instruction tuning stages, we observe that fine-tuning the compressor alongside the generator brings only marginal improvements. For example, under the Normal setting, the average gain of full fine-tuning over generator-only tuning is less than 2.0% across most datasets. Considering that in Section 4.4, the instruction-tuned compressor tends to degrade retrieval performance due to its focus on answer-centric representations, a promising future direction is to explore how to effectively extract task-relevant information from compressed representations without directly fine-tuning the compressor itself. In contrast, during the end-to-end learning stage, fine-tuning the query reasoner proves to be more beneficial. A trainable retrieval module enables the model to identify more relevant documents and provide stronger contextual grounding for the generator. For instance, under the Oracle setting with a compression ratio of 32, the F1 score of Mistral-7B improves from 52.54% to 70.91% when jointly fine-tuning both the query reasoner and generator. This highlights the crucial role of query reasoner in enhancing overall QA performance within our unified training framework.
@@ F.2 Retrieval number generalization
We further explore the impact of varying the number of retrieved documents (top-$k$) during testing in our end-to-end training framework. During training, the model is consistently trained with the top-5 retrieved documents, while at test time, we vary $k$ from 1 to 10 to examine the model’s sensitivity to retrieval size. The results are presented in Figure 7. As shown in the figure, the F1 score generally exhibits a rapid increase followed by a gradual decline as $k$ increases. However, the performance drop remains relatively small, indicating that our trained query reasoner and generator demonstrate good generalization with respect to the number of retrieved documents during inference.
@@ Appendix I Prompts
Figures 916 illustrate the prompts used during the data synthesis process. Specifically, we employ different prompting strategies for (1) generating QA pairs, (2) producing paraphrased documents, (3) validating information completeness, and (4) completing missing information. Additionally, Figure 17 shows the prompt template used by the generation model to answer questions based on the compressed document representations.
@@
Table 21: Instruction-tuning performance with and without the MSE loss under different compression ratios (CR = 32, 128) and normal retrieval settings.

### v3

@@ Appendix E Training Curves
Fig 7 presents the validation curves during end-to-end training on the NQ and Musique datasets. Recall scores consistently increase while evaluation losses steadily decrease, indicating stable and effective optimization. Higher compression ratios generally yield lower recall and higher loss, mirroring the trends observed during the compression pretraining stage.
@@ F.1 Effect of Freezing the Compressor and Query Reasoner
We investigate the effect of limiting the fine-tuning scope to the generator module while freezing both the compressor and query reasoner. Specifically, we examine two representative compression settings, CR=32 and CR=128, and compare model performance when only the generator is fine-tuned during both the instruction tuning and end-to-end QA training stages. The results are shown in Table 12 and 13. During the compression learning and instruction tuning stages, we observe that fine-tuning the compressor alongside the generator brings only marginal improvements. For example, under the Normal setting, the average gain of full fine-tuning over generator-only tuning is less than 2.0% across most datasets. Considering that in Section 4.4, the instruction-tuned compressor tends to degrade retrieval performance due to its focus on answer-centric representations, a promising future direction is to explore how to effectively extract task-relevant information from compressed representations without directly fine-tuning the compressor itself. In contrast, during the end-to-end learning stage, fine-tuning the query reasoner proves to be more beneficial. A trainable retrieval module enables the model to identify more relevant documents and provide stronger contextual grounding for the generator. For instance, under the Oracle setting with a compression ratio of 32, the F1 score of Mistral-7B improves from 52.54% to 70.91% when jointly fine-tuning both the query reasoner and generator. This highlights the crucial role of query reasoner in enhancing overall QA performance within our unified training framework.
@@ F.2 Retrieval number generalization
We further explore the impact of varying the number of retrieved documents (top-$k$) during testing in our end-to-end training framework. During training, the model is consistently trained with the top-5 retrieved documents, while at test time, we vary $k$ from 1 to 10 to examine the model’s sensitivity to retrieval size. The results are presented in Fig 8. As shown in the figure, the F1 score generally exhibits a rapid increase followed by a gradual decline as $k$ increases. However, the performance drop remains relatively small, indicating that our trained query reasoner and generator demonstrate good generalization with respect to the number of retrieved documents during inference.
@@ Appendix I Prompts
Figures 1017 illustrate the prompts used during the data synthesis process. Specifically, we employ different prompting strategies for (1) generating QA pairs, (2) producing paraphrased documents, (3) validating information completeness, and (4) completing missing information. Additionally, Fig 18 shows the prompt template used by the generation model to answer questions based on the compressed document representations.
@@
Table 20: Instruction-tuning performance with and without MSE loss under different compression ratios (CR = 32, 128) and normal retrieval settings.